Two-third of a certain number exceeds a quarter of it by 15.find the 2/3 number?
Answer by Timnewman(323)

let the number be x,
2x/3=x/4+15
2x/3=(x+60)/4
8x=3(x+60)
8x=3x+180
5x=180
x=36
Then
2(36)/3=24
Two third of the number is 24
HOPE THIS HELPS?